Aroldis Chapman posses a 47.7 m/s fastball. If he threw a baseball straight up into the air, how long will the baseball stay in the air?
1
Expert's answer
2020-12-14T12:15:23-0500

The time the baseball will stay in the air is twice its raise time (see http://www.college-physics.com/book/mechanics/vertically-launched-projectiles/). Thus, obtain:



"t = \dfrac{2v_0}{g}"

where "v_2 = 47.7m\/s" is the initial speed, and "g = 9.8m\/s^2" is the gravitational acceleration. Thus, obtain:



"t = \dfrac{2\cdot 47.7m/s}{9.8m/s^2} = 9.7s"

Answer. 9.7s.